OpenStudy (anonymous):

Center in first quadrant, r=3, tangent to y-axis (0,1)?

OpenStudy (raden):

first, graph the circle and utilize after that looks the centre of circle is (3,1)

OpenStudy (raden):

do u want to find an equation of the circle ?

OpenStudy (anonymous):

yes, the question about finding equation of the circle; if it becomes (3.1) then the answer must be like this; (x-3)^2+(y-1)^2=3^2 r2=10 the radius=3 is it correct?

OpenStudy (raden):

the equation of circle be (x-3)^2+(y-1)^2=3^2 u are right ?

OpenStudy (anonymous):

Thank you so much.

OpenStudy (raden):

welcome... but why, u write r^2 = 10 ? it should be r^2 = 9 right ?